You buy an eleven-year bond that has a 8.25% current yield and a 8.25% coupon (paid annually). In one year, promised yields to maturity have risen to 9.25%. What is your holding-period return? homework explanation says:Using a financial calculator, FV = 1,000, n = 10, PMT = 82.50, and i = 9.25 gives us a selling price of $936.52 this year. but when I plugged it in the calculator i get 984.96? please help explain why I'm getting wrong answer in financial calc. answer is 1.9%
Answer #1

Your answer is coming wrong because your calculator is in "BGN" mode

First convert it into "END" mode, you will get correct answer.

You are entering values correct but calculator is the "beginning of the year " mode

you can do this by : pressing "2nd" then pressing "BGN", then again pressing "2nd" and the press "SET"

Now current yield = coupon =8.25%, so bond is trading at par, P0 = purchase price = 1000

After a year, P1 = sale price = 936.52

C1 = annual coupon = 82.50

So Holding period return = (P1 + C1 - P0) / P0 = (936.52 + 82.50 -1000)/ 1000 = 1.9%

Answer : 1.9% (Thumbs up please)